OpenStudy (anonymous):

The density of function X is given by f(x)= a+bx^2 if 0 ≤ x ≤ 1 , 0 otherwise. E(x)=3, find a and b?

OpenStudy (dumbcow):

\[E(x) = \int\limits_{0}^{1}x*f(x) = \int\limits_{0}^{1}ax+bx^{3}\] \[\int\limits\limits_{0}^{1}ax+bx^{3} = 3\] \[\frac{a}{2} + \frac{b}{4} = 3\] \[b = 12-2a\] Now the sum of any probability density function is equal to 1 \[\int\limits_{0}^{1}f(x) = 1\] \[\int\limits_{0}^{1}a+bx^{2} = 1\] \[a+\frac{b}{3} = 1\] \[b=3-3a\] \[\rightarrow 12-2a = 3-3a\] \[a=-9\] \[b=30\]

OpenStudy (zarkon):

dumbcow's solution is correct given the information provided. Unfortunately this gives us a density of \[f(x)=\left\{\begin{matrix}-9+30x^2, & 0≤ x ≤ 1 \\0, & \text{ow}\end{matrix}\right.\] and for small values of \(x\) we have that \(f(x)\) is negative and therefore cannot be a pdf.

OpenStudy (dumbcow):

yeah i noticed that and thought maybe something was off